There are 5 white, 7 red and 8 black balls. we randomly choose 3 ball with replacement . then what is parobability of getting at least one white ball?

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
GivenThere are 5 white, 7 red and 8 black balls
Probability of atleast one white ball is
=p(1)+p(2)+(3)
=(5/20)(15/20)(15/20)+(5/20)(5/20)(15/20)+(5/20)(5/20)(5/20)
=13/64